### v1

@@ Abstract
We study the impact of economic integration on spatial development in a model where all consumers are inter-regionally mobile and have heterogeneous preferences regarding their residential location choices. This heterogeneity is the unique dispersion force in the model. We show that, under reasonable values for the elasticity of substitution among varieties of consumption goods, a higher trade integration always promotes more symmetric patterns, irrespective of the functional form of the dispersion force. We also show that an increase in the degree of heterogeneity in preferences for location leads to less spatial inequality.
@@ 1 Introduction
We show that, irrespective of the degree of heterogeneity, a higher trade integration always promotes more symmetric spatial patterns. Since heterogeneity does not co-vary with transportation costs, the latter only affect the strength of agglomeration forces due to increasing returns. This is a reversion of the usual prediction that lower transport costs lead to agglomeration, and contrasts the findings that consumer heterogeneity leads to a bell-shaped relationship between decreasing transport costs and spatial inequality (Murata, 2003; Tabuchi and Thisse, 2002). Since there is no immobile workforce (no fixed regional internal demand), there is a lower incentive for firms to relocate to smaller markets in order to capture a higher share of local demand and avoid competition when transport costs are higher. Krugman and Elizondo (1996), Helpman (1998) and Murata and Thisse (2005), have also obtained similar results concerning the relationship between transportation costs and spatial inequality. The first includes a congestion cost in the core, the second considers a non-tradable housing sector and the treatment is only numerical, while the latter’s prediction cannot be disassociated from the interplay between inter-regional transportation costs and intra-regional commuting costs. Recently, Tabuchi et al. (2018) have reached similar conclusions to ours arguing that falling transport costs increase the incentives for firms in peripheral regions to increase production since they have a better access to the core, thus contributing to the dispersion of economic activities. Allen and Arkolakis (2014) estimated that the removal of the US Interstate Highway System, which, by limiting accesses, can be interpreted as an increase in transportation costs, would cause a redistribution of the population from more economically remote regions to less remote regions in the US. This adds empirical validity to our findings.
@@ 2 The model
There are two regions $L$ and $R$ with a total population of mass $1$. Consumers are assumed to have heterogeneous preferences with respect to the region in which they reside. We follow Castro et al. (2020) and assume that these preferences are described by a parameter $x,$ uniformly distributed along the interval $[0,1]$. The consumer with preference described by $x=0$ (resp. $x=1$) has the highest preference for residing in region $L$ (resp. region $R$). An agent whose preference corresponds to $x=1/2$ is indifferent between either region. We can thus refer to each consumer with respect to his preference towards living in region $L$ as $x\in[0,1]$.
@@
where $C_{i}$ denotes the level of consumption of a consumer living in region $i\in\left\{L,R\right\}$, and $t(x)$ is the utility penalty associated with living in $L$, while $t(1-x)$ is the utility penalty a consumer faces from living in region $R.$ Regions are symmetric in all respects. We assume that $t^{\prime}(.)>0$. There are two regions $L$ and $R$ with a total population of mass $1$.
The short-run equilibrium is equivalent to Murata (2003) and Tabuchi et al. (2018), which we briefly describe here. The consumption aggregate is a CES composite given by:
@@
$p_{LL}=\dfrac{\beta\sigma}{\sigma-1}w_{L}\text{ \ and\ }p_{RL}=\dfrac{\beta\sigma\tau}{\sigma-1}w_{L}.$ (5)
Labour-market clearing implies that the number of agents in region $L$ equals labour employed by a firm times the number of varieties produced in a region:
@@
from where, using (4), we get:
@@
Figure 1: Short-run equilibrium relative wage as a function of the consumers in region $L$. We set $\sigma=2$ and $\phi=0.5$.
The nominal wage is strictly increasing in the spatial distribution (see Appendix A). Moreover, see Appendix A, when $h>1/2$, the relative wage $w$ increases when trade barriers are higher. Conversely, if $h<1/2$, higher trade barriers implies lower wages in $L$. Therefore higher trade barriers increase the wage divergence between the regions.
The intuition is as follows. When all workers are mobile, they can move to the region that offers them a relatively better access to varieties. This advantage of the larger region becomes higher as transport costs increase because markets become more focused on local demand. This increases expenditures in the larger region relative to the smaller, which in turn pushes the nominal relative wage upwards.
We consider a general isoelastic sub-utility for consumption goods:
$\begin{cases}u_{i}=\dfrac{C_{i}^{1-\theta}-1}{1-\theta},&\text{ if }\theta\in[0,1)\cup(1,+\infty)\\ u_{i}=\ln C_{i},&\text{ if }\theta=1,\end{cases},$ (9)
where for $\theta=1$ we take the limit value of the upper expression of $u_{i}$. The parameter $\theta$ is a positive agglomeration externality. It influences how a change in the regional consumption differential, $C_{L}-C_{R}$, impacts the regional utility differential, $u_{L}-u_{R}$, i.e., it influences the strength of the self-reinforcing agglomeration mechanism when one region is more populated than the other. We shall detail further below why this specification is useful.
$\Delta u=\begin{cases}\dfrac{\eta}{1-\theta}\left\{w^{1-\theta}\left[(1-h)\phi+hw^{1-\sigma}\right]^{\tfrac{1-\theta}{\sigma-1}}-\left[1-h+h\phi w^{1-\sigma}\right]^{\tfrac{1-\theta}{\sigma-1}}\right\},&\ \text{if }\mbox{$\theta\neq$1}\\ \ln w+\dfrac{1}{\sigma-1}\ln\left[\frac{hw^{1-\sigma}+(1-h)\phi}{h\phi w^{1-\sigma}+(1-h)}\right],&\ \text{if $\theta$=1.}\end{cases}$ (10)
We adopt the normalizations by Fujita et al. (1999), i.e., $\alpha\sigma=1$ so that the number of consumers in a region equals its number of firms. Moreover, we assume that $(\sigma-1)/(\sigma\beta)=1$ so that the price of each manufactured variety in a region equals its workers’ nominal wage. These imply that $\eta=1$.
Figure 2 depicts $\Delta u$ in (10) for different levels of $\theta$.
@@
We observe that a higher $\theta$ increases the utility differential $\Delta u$ for any spatial distribution $h>1/2$. Therefore, if $L$ is relatively more industrialized, a higher $\theta$ increases the attractiveness of $L$ relative to $R$ for consumers. Thus, it strengthens the agglomeration forces towards region $L$.
The purpose of the general specification in (9) is explained as follows. Firstly, it allows us to derive general stability conditions for long-run equilibria for any $t(x)$. Secondly, it encompasses the Murata (2003) model as a particular case when $\theta=0$, which is of great interest for comparison purposes.
@@ 3 Long-run equilibria
In a long-run spatial equilibrium, each worker chooses to live in the region that provides a higher utility. Hence, as in Castro et al. (2022), a long-run equilibrium must be such that workers with $x\in[0,h)$ live in region $L$ and workers with $x\in(h,1]$ live in region $R$. Knowing from utility maximization that $C=W/P$ and assuming that $t(x)$ enters additively in overall utility $U_{i}(x)$, we rewrite the indirect utilities as:
@@ 3.1 Interior equilibria
$\zeta\left\{\varphi\left[\dfrac{w^{\sigma}+w^{1-\sigma}-(w+1)\phi}{w\left(1-\phi^{2}\right)}\right]^{\frac{\theta-1}{\sigma-1}}-\psi\left[\dfrac{w^{2\sigma}-(w+1)\phi w^{\sigma}+w}{w\left(1-\phi^{2}\right)}\right]^{\frac{\theta+\sigma-2}{\sigma-1}}\right\}<t^{\prime}\left(h^{*}\right)+t^{\prime}\left(1-h^{*}\right),$
@@
$\zeta$ $=\dfrac{w^{-(\theta+\sigma+1)}}{(\sigma-1)\left[(\sigma-1)\phi+(\sigma-1)\phi w^{2\sigma}+\left(-2\sigma+\phi^{2}+1\right)w^{\sigma}\right]};$
$\varphi$ $=w\left\{\phi\left[\sigma+(\sigma-1)w\right]w^{\sigma}-2\sigma w+w\right\}\left[w^{2\sigma}-(w+1)\phi w^{\sigma}+w\right];$
$\psi$ $=\left(1-\phi^{2}\right)w^{\theta+\sigma+1}\left[(2\sigma-1)w^{\sigma}-\sigma(w+1)\phi+\phi\right].$
@@
A careful inspection of (12) allows us to conclude that the LHS is decreasing in $\phi$ if $\theta\geq 1$. For $\theta<1$, it is decreasing in $\phi$ if $\sigma\geq 1+\sqrt{2}/2\approx 1.71$. Symmetric dispersion then becomes easier to sustain under lower transportation costs if $\sigma>1+\sqrt{2}/2$, which, according to recent empirical estimations for $\sigma$, is more than reasonable. [^3] Estimations evidence that $\sigma$ should be significantly larger than unity (Crozet, 2004; Head and Mayer, 2004; Niebuhr, 2006; Bosker et al., 2010). Anderson and Wincoop (2004), for instance, find that it is likely to range from 5 to 10. Since $t(h)$ does not depend on $\phi$, this result does not depend on consumer heterogeneity. Since, at symmetric dispersion, consumers in each region have access to the same amount of manufactures, an exogenous migration will induce a lower (higher) benefit from local consumption goods in the larger market if transport costs are lower (higher). This is captured by the fact that the relative decrease in prices and increase in wages (at the symmetric equilibrium) is more pronounced when transport costs are higher.
Let us now assume that a partial agglomeration $h^{*}\in(1/2,1)$ equilibrium exists and is unique. Regarding the influence of trade integration on partial agglomeration equilibria, we establish the following result.
@@
If a stable partial agglomeration $h^{*}\in(1/2,1)$ exists and is unique, it becomes more symmetric as trade barriers decrease.

### v2

@@ Abstract
We study the impact of economic integration on agglomeration in a model where all consumers are inter-regionally mobile and have heterogeneous preferences regarding their residential location choices. This heterogeneity is the unique dispersion force in the model. We show that, under reasonable values for the elasticity of substitution among varieties of consumption goods, a higher trade integration always promotes more symmetric spatial patterns, reducing the spatial inequality between regions, irrespective of the functional form of the dispersion force. We also show that an increase in the degree of heterogeneity in preferences for location leads to a more even spatial distribution of economic activities and thus also reduces the spatial inequality between regions.
@@ 1 Introduction
We show that, irrespective of the degree of heterogeneity, a higher trade integration always promotes a more symmetric spatial distribution of economic activities, thus reducing the spatial inequality in terms of industry size between regions. Since heterogeneity does not co-vary with transportation costs, the latter only affect the strength of agglomeration forces due to increasing returns. This is a reversion of the usual prediction that lower transport costs lead to agglomeration (Krugman, 1991; Fujita et al., 1999, Chap. 5), and contrasts the findings that consumer heterogeneity leads to a bell-shaped relationship between decreasing transport costs and the spatial distribution of economic activities (Murata, 2003; Tabuchi and Thisse, 2002). Since there is no immobile workforce (no fixed regional internal demand), there is a lower incentive for firms to relocate to smaller markets in order to capture a higher share of local demand and avoid competition when transport costs are higher. Krugman and Elizondo (1996), Helpman (1998) and Murata and Thisse (2005), have also obtained similar results concerning the relationship between transportation costs and the spatial distribution of economic activities. The first includes a congestion cost in the core, the second considers a non-tradable housing sector and the treatment is only numerical, while the latter’s prediction cannot be disassociated from the interplay between inter-regional transportation costs and intra-regional commuting costs. Recently, Tabuchi et al. (2018) have reached similar conclusions to ours arguing that falling transport costs increase the incentives for firms in peripheral regions to increase production since they have a better access to the core, thus contributing to the dispersion of economic activities. Allen and Arkolakis (2014) estimated that the removal of the US Interstate Highway System, which, by limiting accesses, can be interpreted as an increase in transportation costs, would cause a redistribution of the population from more economically remote regions to less remote regions in the US. This adds empirical validity to our findings.
@@ 2 The model
There are two regions $L$ and $R$, symmetric in all respects, with a total population of mass $1$. Consumers are assumed to have heterogeneous preferences with respect to the region in which they reside. We follow Castro et al. (2022) and assume that these preferences are described by a parameter $x,$ uniformly distributed along the interval $[0,1]$. The consumer with preference described by $x=0$ (resp. $x=1$) has the highest preference for residing in region $L$ (resp. region $R$). An agent whose preference corresponds to $x=1/2$ is indifferent between either region. We can thus refer to each consumer with respect to his preference towards living in region $L$ as $x\in[0,1]$.
@@
where $C_{i}$ denotes the level of consumption of a consumer living in region $i\in\left\{L,R\right\}$, and $t:[0,1]\mapsto\mathbb{R}$ is the utility penalty of consumer $x$ associated with living in $L$, while $t(1-x)$ is the utility penalty the same consumer faces from living in region $R.$ We assume that $t(x)$ is differentiable, such that $t^{\prime}(x)>0,\forall x\in[0,1]$.
The number of agents in region $L$ is given by $h\in[0,1]$ and fully describes the spatial distribution of economic activities. In the short-run, the spatial distribution of workers and firms is fixed and an equilibrium is reached when all prices, wages, and output adjust to clear markets given $h$. In what follows, the short-run equilibrium is equivalent to Murata (2003) and Tabuchi et al. (2018), which we briefly describe here. The consumption aggregate is a CES composite given by:
@@
$p_{LL}=\dfrac{\beta\sigma}{\sigma-1}w_{L}\text{ \ and\ }p_{RL}=\dfrac{\beta\sigma\tau}{\sigma-1}w_{L},$ (5)
where $p_{ij}$ is the price of a good produced in region $i$ and consumed in region $j$. Labour-market clearing implies that the number of agents in region $L$, equals labour employed by a firm times the number of varieties (and hence firms) $n_{L}$ produced in region $L$:
@@
from where, using (4), we get the number of firms in each region $i$:
@@
where $\phi\equiv\tau^{1-\sigma}\in(0,1)$ is the freeness of trade.
The nominal wage $w$ can be implicitly defined as a function of the spatial distribution in region $L$, $h\in[0,1]$. [^1] The conditions for application of the Implicit Function Theorem are shown to be satisfied in Appendix A. We have $w(0)=\phi^{1/\sigma}<1$, $w(1/2)=1$ and $w(1)=\phi^{-1/\sigma}>1$. We say that $L$ is larger than $R$ when $h>1/2,$ and smaller otherwise.
Figure 1: Short-run equilibrium relative wage as a function of the consumers in region $L$. We set $\sigma=2$ and $\phi\in\{0.1,0.5,0.7\}$.
Figure 1 illustrates $w(h)$ for $h\in[0,1]$ with parameter values $\sigma=2$ and $\phi\in\{0.1,0.5,0.7\}$. The following result corroborates this illustration and discusses the impact of trade barriers on wages.
Lemma 1.
The relative nominal wage $w$ is strictly increasing in $h$. When region $L$ is larger (smaller) than region $R$, the wage $w$ is decreasing (increasing) in $\phi$.
Proof.
See Appendix A. ∎
We conclude that higher trade barriers increase the wage divergence between the regions. The intuition is as follows. When all workers are mobile, they can move to the region that offers them a relatively better access to varieties. This advantage of the larger region becomes higher as transport costs increase because markets become more focused on local demand. This increases expenditures in the larger region relative to the smaller, which in turn pushes the nominal relative wage upwards.
We consider a general isoelastic sub-utility for consumption goods:
$\begin{cases}u_{i}=\dfrac{C_{i}^{1-\theta}-1}{1-\theta},&\text{ if }\theta\in[0,1)\cup(1,+\infty)\\ u_{i}=\ln(C_{i}),&\text{ if }\theta=1,\end{cases},$ (9)
where for $\theta=1$ we take the limit value of the upper expression of $u_{i}$. The parameter $\theta$ is a positive agglomeration externality. It influences how a change in the regional consumption differential, $C_{L}-C_{R}$, impacts the regional utility differential, $u_{L}-u_{R}$, i.e., it influences the strength of the self-reinforcing agglomeration mechanism when one region is more populated than the other. The specification in (9) is quite general and encompasses the Murata (2003) model as a particular case when $\theta=0$, which shall be of great interest for comparison purposes in Section 4.222We note in advance that the qualitative results of the model obtained in subsequent sections do not depend on $\theta$ (c.f., Section 3.2).
$\Delta u=\begin{cases}\dfrac{\eta}{1-\theta}\left\{w^{1-\theta}\left[(1-h)\phi+hw^{1-\sigma}\right]^{\tfrac{1-\theta}{\sigma-1}}-\left[1-h+h\phi w^{1-\sigma}\right]^{\tfrac{1-\theta}{\sigma-1}}\right\},&\ \text{if }\mbox{$\theta\neq 1$}\\ \ln w+\dfrac{1}{\sigma-1}\ln\left[\frac{hw^{1-\sigma}+(1-h)\phi}{h\phi w^{1-\sigma}+(1-h)}\right],&\ \text{if }\theta=1.\end{cases}$ (10)
@@
Figure 2 depicts $\Delta u$ in (10) for different levels of $\theta$. We observe that a higher $\theta$ increases the utility differential $\Delta u$ for any spatial distribution $h>1/2$. Therefore, if $L$ is relatively more industrialized, a higher $\theta$ increases the attractiveness of $L$ relative to $R$ for consumers. Thus, it strengthens the agglomeration forces towards region $L$.
@@ 3 Long-run equilibria
In a long-run spatial equilibrium, each worker chooses to live in the region that provides a higher utility. We follow Castro et al. (2022) and assume that $t(x)$ enters additively in overall utility, $U_{i}(x)$, so that the utility penalty, and hence heterogeneity in preferences for location, are modelled à la Hotelling (1929). [^4] It could be interesting to consider the case where $t(x)$ is multiplicative (e.g. as a percentage of $U_{i}(x)$). However, the non-linear impact of $t(x)$ would likely make analytical results harder to obtain. In a long-run equilibrium, workers with $x\in[0,h)$ live in region $L$, and workers with $x\in(h,1]$ live in region $R$. Knowing from utility maximization that $C_{L}=w/P_{L}$ and $C_{R}=1/P_{R}$, we rewrite the indirect utilities of the consumer with $x=h$ as:
@@ 3.1 Interior equilibria
$\zeta\left\{\left(\varphi w^{\sigma\frac{1-\theta}{\sigma-1}}+\frac{\psi}{w}\right)\left\{\frac{\left(1-\phi^{2}\right)w}{w^{2\sigma}-\left[\varphi(w+1)w^{\sigma}\right]+w}\right\}^{\frac{1-\theta}{\sigma-1}}\right\}<t^{\prime}\left(h^{*}\right)-t^{\prime}\left(1-h^{*}\right),$
@@
$\zeta$ $=\dfrac{w^{2\sigma}-\left[\phi(w+1)w^{\sigma}\right]+w}{(\sigma-1)\left[(\sigma-1)\phi+(\sigma-1)\phi w^{2\sigma}+\left(-2\sigma+\phi^{2}+1\right)w^{\sigma}\right]};$
$\varphi$ $=w^{-\theta-\sigma}\left\{\phi\left[\sigma+(\sigma-1)w\right]w^{\sigma}-2\sigma w+w\right\};$
$\psi$ $=(\sigma-1)\phi+(1-2\sigma)w^{\sigma}+\sigma\phi w.$
@@
As expected, symmetric dispersion is stable if, after a small increase in $h$, the relative utility gain from consumption is lower than the increase in the utility penalty for the agent of type $x=h$. A careful inspection of (12) allows us to conclude that the LHS is decreasing in $\phi$ if $\theta\geq 1$. For $\theta<1$, it is decreasing in $\phi$ if $\sigma\geq 1+\sqrt{2}/2\approx 1.71$. Symmetric dispersion then becomes easier to sustain under lower transportation costs if $\sigma>1+\sqrt{2}/2$, which, according to recent empirical estimations for $\sigma$, is more than reasonable. [^5] Estimations evidence that $\sigma$ should be significantly larger than unity (Crozet, 2004; Head and Mayer, 2004; Niebuhr, 2006; Bosker et al., 2010). Anderson and Wincoop (2004), for instance, find that it is likely to range from 5 to 10. As $\phi\rightarrow 1$, symmetric dispersion is always stable because $t^{\prime}\left(\tfrac{1}{2}\right)>0$. This warrants the following result.
@@
As $\phi$ increases from a low level, symmetric dispersion either remains stable or becomes stable, if either conditions hold: (i). $\theta\in[0,1)$ and $\sigma>1+\sqrt{2}/2$; or (ii). $\theta\geq 1$.
Thus, an increase in $\phi$ can never turn symmetric dispersion from stable to unstable, suggesting that lower trade barriers always promote an even distribution of economic activities. Since, at symmetric dispersion, consumers in each region have access to the same amount of manufactures, an exogenous migration will induce a lower (higher) benefit from local consumption goods in the larger market if transport costs are lower (higher). This is captured by the fact that the relative decrease in prices and increase in wages (at the symmetric equilibrium) is more pronounced when transport costs are higher.
Let $\phi=\phi_{b}\in(0,1)$ such that $\frac{d\Delta u}{dh}\left(\tfrac{1}{2};\phi_{b}\right)=t^{\prime}\left(\tfrac{1}{2};\phi_{b}\right)$. If $\phi_{b}$ exists, then symmetric dispersion is stable for $\phi>\phi_{b}$ and unstable for $\phi<\phi_{b}$. We have the following result.
Lemma 3.
A curve of partial agglomeration equilibria $h^{*}\in(1/2,1)$ branches from $\phi=\phi_{b}.$ If it branches in the direction of $\phi<\phi_{b}$, partial agglomeration is stable.
Proof.
See Appendix B. ∎
The result above establishes existence of partial agglomeration in a neighbourhood of $\phi=\phi_{b}$. [^6] Uniqueness is not ensured for the entire range of $\phi\in(0,\phi_{b})$ as the model may undergo secondary bifurcations along the primary branch of partial agglomeration equilbria, such as a saddle-node bifurcation (see e.g. Gaspar et al. (2018, 2021) and Castro et al. (2021) for applications in economic geography). Let us now assume that a partial agglomeration $h^{*}\in(1/2,1)$ equilibrium exists and is stable for some value of $\phi\in(0,1)$. Regarding the influence of trade integration on partial agglomeration equilibria, we establish the following result.
Proposition 2.
@@
If a stable partial agglomeration $h^{*}\in(1/2,1)$ exists, it becomes more symmetric as trade barriers decrease.
Proof.
See Appendix B. ∎
This result states that if most consumers reside in region $L$, increasing the freeness of trade will lead to a smooth exodus from region $L$ to region $R$, irrespective of the value of $t(h^{*})$. We conclude from Propositions 1 and 2 that more integration leads agents to distribute more equally among the two regions, as in Helpman (1998), who also considers that all labour is inter-regionally mobile. This qualitative effect of $\phi$ does not depend on consumer heterogeneity. It thus contrasts the findings that consumer heterogeneity leads to a bell-shaped relationship between decreasing transport costs and agglomeration (Murata, 2003; Tabuchi and Thisse, 2002).
3.2 Logit preferences for location
We begin this Section by noting that, for any value of $\theta$, there exists a non-trivial subset of a suitably defined parameter space over which the qualitative results of the model remain invariant. Therefore, for the remainder of this Section, we will assume that utility is logarithmic in consumption. This greatly simplifies the analysis and entails no loss of generality.
Assumption 1.
Utility $u_{i}$ is logarithmic in $C_{i}$: $\theta=1\implies u_{i}=\ln(C_{i})$.
For the two region case, and following Tabuchi and Thisse (2002) and Murata (2003), we assume that the probability that a consumer will choose to reside in region $L$ is given by the Logit model, written as:
$\Phi_{L}(h)=\dfrac{e^{u_{L}(h)/\mu}}{e^{u_{L}(h)/\mu}+e^{u_{R}(h)/\mu}},$ (13)
where $\mu\geq 0$ is a scale parameter which measures the dispersion of consumer preferences. If $\mu=0$, consumers do not care about their location preferences but rather solely about relative wages. The number of agents $h$ is the same as the consumer $x$ who is indifferent between living in region $L$ or in region $R$. Therefore, the probability $\Phi_{L}(h)$ is tantamount to the indifferent consumer $x=h$ (as in Castro et al., 2022). Thus, using (13), we can write:
$h=\dfrac{e^{u_{L}(h)/\mu}}{e^{u_{L}(h)/\mu}+e^{u_{R}(h)/\mu}}.$ (14)
Manipulating (14) yields:
$u_{L}+\mu\ln(1-h)=u_{R}+\mu\ln(h),$ (15)
which rearranged yields the long-run equilibrium condition, $\Delta u(h)=\Delta t(h),$ with $t(x)=-\mu\ln(1-x).$ This yields $\Delta t(h)=\mu\left[\ln h-\ln(1-h)\right]$, dubbed by Castro et al. (2022) as the “home-sweet-home” effect, which comprises the model’s unique dispersion force generated by consumer heterogeneity in preferences for location.
Notice from (15) that $t(h)=-\mu\ln(1-h)>0$ for $h\in[0,1]$. The overall utility of a consumer in region $i$ thus lies on the interval $(-\infty,u_{i}]$. For a strictly positive $\mu$, the consumer who likes region $R$ the most (at $x=1$) will never want to live in $L$, because, at $h=1$, the overall utility loss from living in region $L$ is unbounded. This means that the Logit model penalizes (benefits) the consumers who are less (more) willing to leave a region very strongly. Therefore, agglomeration in any region ($h\in\{0,1\})$ is not possible. Now, suppose condition (12) does not hold and symmetric dispersion is unstable. Then, we have $\frac{d\Delta u}{dh}\left(\tfrac{1}{2}\right)>t^{\prime}\left(\tfrac{1}{2}\right)$ which means that $\Delta u(h)>t(h)-t(h)$ for $h=\tfrac{1}{2}+\epsilon,$ with $\epsilon>0$ small enough. Since $\Delta u\left(1\right)-\Delta t(1)\rightarrow-\infty$, then, by the Intermediate Value Theorem, we must have $\Delta V=0$ for some $h\in(1/2,1)$. Accordingly, a partial agglomeration equilibrium $h^{*}\in(1/2,1)$ must always exist when symmetric dispersion is unstable.
The discussion above allows us to formalize our next result, which summarizes the possible spatial outcomes under Logit type preferences, depending on the degree of heterogeneity $\mu.$
Proposition 3.
Under Logit type preferences and Assumption 1, the spatial distribution depends on the level of consumer heterogeneity as follows:
  • Symmetric dispersion is the only stable equilibrium if:
    $\mu>\mu_{d}\equiv\frac{(2\sigma-1)(1-\phi)}{(\sigma-1)(2\sigma+\phi-1)};$ (16)
  • Partial agglomeration $h^{*}\in\left(\tfrac{1}{2},1\right)$ is the only stable equilibrium and is unique if $\mu<\mu_{d}$.